aliviya06:

Andy flew his model airplane at full speed straight into the wind for 10 seconds before turning it around and flying with the wind for 5 seconds, again at full speed. The plane flew 50 yards on the way out and 45 yards on the way back. How fast was the wind blowing that day in yd/sec?

martai:

what are the answer choices ?

martai:

??

aliviya06:

A. 2 yd/sec B. 5 yd/sec C. 6 yd/sec D. 7 yd/sec

martai:

oki

Vocaloid:

in general, speed = distance/time let's let a = the speed of the airplane by itself, with no wind let w = the speed of the wind when he flies the airplane into the wind, the wind slows down the airplane. the speed of the airplane is (a - w). the distance is 50 and the time is 10, giving us (a-w) = 50/10 = 5 for the return trip, he's flying with the wind, so the wind speeds up the airplane. the speed of the airplane on the way back is (a + w). the distance is 45 and the time is 5 giving us: (a+w) = 45/5 = 9 we want to solve for w, the speed of the wind. subtract the two equations to eliminate a, solve for w.